OpenStudy (anonymous):

Hellllo, I need some help with Math. Please help. Hellllo, I need some help with Math. Please help. @Mathematics

OpenStudy (anonymous):

Here is the file.

OpenStudy (anonymous):

Any line can be represented as y = mx + b where m is the slope and b is the y-intercept (where x = 0). You are given two points. Slope is "change in y" divided by "change in x". Calculate slope using that formula given the two points.

OpenStudy (anonymous):

I found the slope to be 2/3 How do you place it in standard form?

OpenStudy (anonymous):

Once you find slope. Find B by substituting one of the points into the y=mx+b equation.

OpenStudy (anonymous):

"b" I mean. Not B.

OpenStudy (anonymous):

Ok so here it is y=2/3 x + 4

hero (hero):

y - 3 = 2/3(x+2) simplify that

OpenStudy (anonymous):

Cool! Then, you simply multiple 3 through. You get 3y = 2x + 12.

OpenStudy (anonymous):

I need it in Ax+By=C form

OpenStudy (anonymous):

multiply. not multiple.

OpenStudy (anonymous):

Sure. If 3y = 2x + 12, you can re-write it as 2x - 3y = -12.

OpenStudy (anonymous):

I need it in Standard form

OpenStudy (anonymous):

Just a bunch of re-arrangement. You did awesome hard part in getting the equation right in the first place. Impressive!

OpenStudy (anonymous):

Thanks a lot man. I really appreciate that. I have had some trouble with this problem, and I had posted it before, but no one was able to help me. Here is the question: Input in standard form the equation of the given line.The line that passes through (-2, 4) and is parallel to x - 2 y = 6

OpenStudy (anonymous):

Do you think you can help me?

OpenStudy (anonymous):

Yes. Step at a time.

OpenStudy (anonymous):

Yeah that is good.

OpenStudy (anonymous):

First convert to y=mx+b form. Makes life easy. You can re-arrange at the end.

OpenStudy (anonymous):

x-2y=6 written in y=mx+b form is?

OpenStudy (anonymous):

2y=-x+6

OpenStudy (anonymous):

I don't think that is right.

OpenStudy (anonymous):

It is mostly there...now take 2 to the other side and divide.

OpenStudy (anonymous):

y=1/2x+3

OpenStudy (anonymous):

x - 2y = 6 => - 2y = -x + 6 => y = 1/2x - 3

OpenStudy (anonymous):

Oh I got the last part wrong.

OpenStudy (anonymous):

No worries. Now, we have it in the form of y = mx + b where m is the slope. So, in this case slope is?

hero (hero):

I thought I was supposed to be helping you. Who is this other guy? lol

OpenStudy (anonymous):

Well you can help me with another lol...

OpenStudy (anonymous):

1/2x

OpenStudy (anonymous):

Well, slope is 1/2. Not 1/2x.

OpenStudy (anonymous):

Oh haha

OpenStudy (anonymous):

The problem says the second line is parallel to this one. The one thing about parallel lines is that they all have the sale slope. So, if the second line is of the form y = mx +b, then it has to be y = 1/2x+b.

OpenStudy (anonymous):

Ok I guess that makes sense.

OpenStudy (anonymous):

It also says that the second line passes through the point (-2, 4). Plugging in -2 for x and 4 for y, you should be able to find b of the second line and then re-write it in standard form.

OpenStudy (anonymous):

Can you speak in English please?

OpenStudy (anonymous):

y = 1/2x + b is our second line. This line passes through the point (-2,4).

OpenStudy (anonymous):

Point (-2,4) simply means x = -2 and y = 4 at that point.

OpenStudy (anonymous):

4y=1/2(-2) *b

OpenStudy (anonymous):

4 = 1/2(-2) + b

OpenStudy (anonymous):

4=-1+b

OpenStudy (anonymous):

Yes. So b=?

OpenStudy (anonymous):

Yeah I don't know was the the other one.

OpenStudy (anonymous):

b = 4+1 = 5. So, the line must be y = 1/2x + 5.

OpenStudy (anonymous):

In standard form... it is 1/2x-y=5

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Multiply by 2 through. 2y = x + 10 x - 2y = -10

OpenStudy (anonymous):

Sorry, your way is fine too. Just more clean when there are no fractions involved.

OpenStudy (anonymous):

Oh yeah I forgot about that step.

OpenStudy (anonymous):

Yup that is Ok... This one should be easy but I do't think it is. Input in standard form the equation of the given line.The line with m = 1/2 and b = 3.

OpenStudy (anonymous):

We already have the slope and the b

OpenStudy (anonymous):

y = mx + b where m = 1/2 and b = 3.

OpenStudy (anonymous):

6x+2y=6

OpenStudy (anonymous):

2x=2y=6

OpenStudy (anonymous):

That is correct right

OpenStudy (anonymous):

y = 1/2x + 3 multiplying by 2 through, you get 2y = x + 6 x - 2y = -6